QUESTION 2 Find lim xy +1 (x, y)*(0,0) x2 +y2-1 O-1 O 2 00 O Does not exist.
Answer #1


\lim_{\left (x,y \right )\rightarrow \left ( 0,0 \right )}\frac{xy+1}{x^{2}+y^{2}-1}

             =\frac{0+1}{0^{2}+0^{2}-1}

               =-1

\lim_{\left (0,y \right )\rightarrow \left ( 0,0 \right )}\frac{0+1}{0+y^{2}+-1}=\frac{1}{0-1}=-1

\lim_{\left (x,0\right )\rightarrow \left ( 0,0 \right )}\frac{0+1}{x^{2}+0+-1}=\frac{1}{0-1}=-1

\lim_{\left (x,y \right )\rightarrow \left ( 0,0 \right )}\frac{xy+1}{x^{2}+y^{2}-1}=-1

Answer : -1